Using your calculator, run a regression analysis on the following bivariate set of data with y as the response variable.

x y

71.2 23.1

90.8 122.9

88.8 82.9

57.7 18.3

71.4 8.6

60.4 25.9

60.2 -43.1

88.5 77.4

68.2 -5.9

41.1 -83

61.4 -2.5

87.1 39.4

Find the correlation coefficient and report it accurate to three decimal places. r =

What proportion of the variation in y can be explained by the variation in the values of x? Report answer as a percentage accurate to one decimal place. (If the answer is 0.84471, then it would be 84.5%...you would enter 84.5 without the percent symbol.) r² = %

Based on the data, calculate the regression line (each value to three decimal places) y = x +

Predict what value (on average) for the response variable will be obtained from a value of 82.2 as the explanatory variable. Use a significance level of α = 0.05 to assess the strength of the linear correlation.

What is the predicted response value? (Report answer accurate to one decimal place.) y =
Answer #1

533.61 5069.44 23.1 1644.72 71.2 1 8244.64 15104.41 11159.32 122.9 90.8 6872.41 7885.44 7361.52 82.9 88.8 334.89 1055.91 3329The correlation coefficient r is computed using the following expression SSxy where Ху. 7l 12 In this case, based on the dataTherefore, based on this information, the sample correlation coefficient is computed as follows 81011 V2672.187 x 34112.88 SS

Proportion of the variation in y can be explained by the variation in the values of x = r2 = 0.8892 = 79.0%

Now,

Therefore, based on the above calculations, the regression coefficients (the slope m, and the y- intercept n) are obtained as

The regression line is : - 202.205 + 3.177 X

Predicted response value = - 202.205 + 3.177*82.2 = 58.9

Significance of correlation coefficient :

The following needs to be tested:

Ho : ρ=0

H_1: \rho\neq0

where ρ corresponds to the population correlation 12, so then the number of degrees of freedom is df = n-2-12-2-10 The sample

Hence, Linear correlation is significant at 5% level of significance.